Trains A and B are traveling in the same direction on parallel tracks. Train A is traveling at 60 mph. Train B is traveling at 64 mph. Train A passes a station at 1:15pm. If Train B passes the same station at 1:45 pm at what time will train B catch up to Train A?
Answer by geetha_rama(94)

Speed of train A = 60mph
Speed of train B = 64mph
At 1:15PM Train A passes a station and train B passes the same station at 1:45PM.
Distance travelled by Train A from 1:15 pm to 1:45 pm = 60*1/2 = 30 miles.
Assume that it takes t hours from 1:45 PM for train B to catch train A
Let Distance travelled by train A in t hours = x miles - > equation 1
Distance that train B should travel in t hours = x+ 30 miles -> equation 2( since train A is already 30miles ahead of train B)
from equaion 1,
x = 60t
from equation 2,
x+30 = 64t
substituting value of x,
60t+ 30 = 64t
4t = 30
=> t = 7.5
Train B will take 7 and a half hours to cross train B
At 9:15PM train B will cross train A (1:45 + 7:30)